Diogo has a utility function U(B, Z) = ABαZβ, where A, α, and β are constants, B is burritos, and Z is pizzas. If the price of burritos, pB, is $2 and the price of pizzas, pZ, is $1, and Y is $100, what is Diogo’s optimal bundle?


1
Expert's answer
2021-06-22T10:18:05-0400

If Diogo has a utility function"U(B, Z) = AB^{\alpha} Z^{\beta};" and if the price of burritos, Pb, is N$2 and the price of pizzas, Pz is N$1, and Y is N$100, using the Lagrangian method we can find the optimal bundle:


"Qb = \frac{\beta}{(\alpha+ \beta)}\times\frac{100}{2} = \frac{50\beta}{(\alpha + \beta)}"


"Qz = \frac{100\beta}{(\alpha + \beta)}"